# Subvarieties of pointed Abelian $\ell$-groups

### v1

@@ 5 Axiomatization of subvarieties of $p\mathbb{AL}$
Now let us fix $p\in I\setminus J$. Since $p\notin J$ by the same argument as in the previous paragraph we obtain ${\boldsymbol{\mathit{Z}}}_{j}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0}\vDash\mathrm{(div_{p,n})}$ for each $j\in J$. By Lemma 5.1 we have ${\boldsymbol{\mathit{Z}}}_{i}\vDash\mathrm{(\text{s-rank}_{n})}$ for each $i\in I$ and thus by Lemma 5.5 we get ${\boldsymbol{\mathit{Z}}}_{i}\vDash\mathrm{(mix_{p,n})}$ and ${\boldsymbol{\mathit{Z}}}_{j}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0}\vDash\mathrm{(mix_{p,n})}$ for each $i\in I$ and $j\in J$. Thus $V_{I,J}\vDash\mathrm{(mix_{p,n})}$ for each $p\in I\setminus J$.
@@
This concludes that $V_{I,J}\vDash\mathrm{(eq)}$ for each $\mathrm{(eq)}\in S_{I,J}$. It remains to show $V_{I,J}$ is uniquely determined by the equations from $S_{I,J}$.
@@
Let us have a finitely generated ${\boldsymbol{\mathit{A}}}_{a}\in p\mathbb{AL}^{+}_{SI}$ satisfying all the equations from $S_{I,J}$. We will show ${\boldsymbol{\mathit{A}}}_{a}\in V_{I,J}$. First, let us note that clearly $\mathbf{HSP}({\boldsymbol{\mathit{A}}}_{a})\neq\mathbf{HSP}({\boldsymbol{\mathit{R}}}_{1})$ since from Lemma 3.7 we know $\mathbf{HSP}({\boldsymbol{\mathit{R}}}_{1})=p\mathbb{AL}^{+}$ and $\boldsymbol{\mathit{A}}$} satisfies some equations, which are non-trivial in $p\mathbb{AL}^{+}$. By Theorems 3.10 and 4.10 it’s enough to consider cases ${\boldsymbol{\mathit{A}}}_{a}\in\{{\boldsymbol{\mathit{Z}}}_{i},{\boldsymbol{\mathit{Z}}}_{j}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0}\mid i,j\in\mathbb{N}\}$.
@@
For ${\boldsymbol{\mathit{A}}}_{a}={\boldsymbol{\mathit{Z}}}_{m}$ we get $i\leq n$, by applying Lemma 5.1 using ${\boldsymbol{\mathit{Z}}}_{m}\vDash\mathrm{(rank_{n})}$. Since ${\boldsymbol{\mathit{Z}}}_{m}\vDash\mathrm{(div_{p,n})}$ for all $p\notin I$ we get by Lemma 5.2 that $m\in I$.
@@
For ${\boldsymbol{\mathit{A}}}_{a}={\boldsymbol{\mathit{Z}}}_{m}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0}$ we similarly get $m\leq n$ and $m\in I$. Since for every $p\in I\setminus J$ we have ${\boldsymbol{\mathit{Z}}}_{m}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0}\vDash\mathrm{(mix_{p,n})}$, by Lemma 5.5 we get $m\in J$.
Thus ${\boldsymbol{\mathit{A}}}_{a}\in V_{I,J}$, which completes the proof. ∎
@@
The structural part of Theorem 5.7 was proved in [28]. This result can be obtained by applying the Mundici functor, Theorem 3.10, and Komori classification, an approach we will discuss later in Section 6 and demonstrate in Theorems 6.3 and 6.6.
@@
Despite the fact that the statement of Theorem 5.7 was about proper varieties, we can also claim that any variety of $p\mathbb{AL}^{+}$ is equal to $V_{I,J}$ for some sets $J\subseteq I$, since $p\mathbb{AL}^{+}=V_{\mathbb{N},\mathbb{N}}$.
@@
Any proper subvariety of $p\mathbb{AL}^{-}$ is of the form
$V^{d}_{I,J}=\mathbf{HSP}({\boldsymbol{\mathit{Z}}}_{i},{\boldsymbol{\mathit{Z}}}_{j}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0}\mid i\in I,j\in J)$
for some finite sets $J\subseteq I$.
Moreover, $V^{d}_{I,J}$ is axiomatized by the following set $S_{I,J}^{d}$ of equations:
@@
Any subvariety of $p\mathbb{AL}$ can be written as ${{\mathbb{K}}}=V_{I_{1},J_{1}}\lor V^{d}_{I_{2},J_{2}}$ for some (not necessarily finite) sets $I_{1},I_{2},J_{1},J_{2}$. Such a variety is axiomatized by the equations from $S_{I_{1},J_{1}}\cup S_{I_{2},J_{2}}^{d}$, where $S_{I_{1},J_{1}}$ and $S^{d}_{I_{2},J_{2}}$ are defined in Theorems 5.7 and 5.8.
@@
Let $\mathbb{K}$}} be arbitrary subvariety of $p\mathbb{AL}$. Using Theorem 2.5 we can derive that subdirectly irreducible pointed Abelian $\ell$-groups are totally ordered. Let us denote ${{\mathbb{K}}}_{SI}^{+}={{\mathbb{K}}}_{SI}\cap p\mathbb{AL}^{+}$ and ${{\mathbb{K}}}_{SI}^{-}={{\mathbb{K}}}_{SI}\cap p\mathbb{AL}^{-}$. Thus we have
$\mathbf{HSP}({{\mathbb{K}}}_{SI}^{+}\cup{{\mathbb{K}}}_{SI}^{-})=\mathbf{HSP}({{\mathbb{K}}}_{SI})={{\mathbb{K}}}.$
This shows that $\mathbb{K}$}} is the join of $\mathbf{HSP}({{\mathbb{K}}}_{SI}^{+})$ and $\mathbf{HSP}({{\mathbb{K}}}_{SI}^{-})$. By Theorem 5.7 $\mathbf{HSP}({{\mathbb{K}}}_{SI}^{+})=V_{I_{1},J_{1}}$ for some sets $I_{1},J_{1}$ and by Theorem 5.8 $\mathbf{HSP}({{\mathbb{K}}}_{SI}^{-})=V^{d}_{I_{2},J_{2}}$ for some sets $I_{2},J_{2}$. Thus we have ${{\mathbb{K}}}=V_{I_{1},J_{1}}\lor V^{d}_{I_{2},J_{2}}$.
We show that the equations from $S_{I_{1},J_{1}}\cup S_{I_{2},J_{2}}^{d}$ axiomatize the variety $\mathbb{K}$}}. First all the equations from $S_{I_{1},J_{1}}$ are valid in all negative pointed Abelian $\ell$-group by Lemmas 5.2, 5.4 and 5.5. Similarly, all the equations from $S^{d}_{I_{2},J_{2}}$ are valid in all positively pointed Abelian $\ell$-groups. Since all the equations from $S_{I_{1},J_{1}}$ are valid in $V_{I_{1},J_{1}}$ and all the equations from $S^{d}_{I_{2},J_{2}}$ are valid in $V^{d}_{I_{1},J_{1}}$ we conclude that all equations from $S_{I_{1},J_{1}}\cup S_{I_{2},J_{2}}^{d}$ are valid in $\mathbb{K}$}}.
Now let us take arbitrary ${\boldsymbol{\mathit{A}}}_{a}\in p\mathbb{AL}_{SI}$. By Theorem 2.5 ${\boldsymbol{\mathit{A}}}_{a}$ is totally ordered. We distinguish three cases:
@@
    If $a=0$ then ${\boldsymbol{\mathit{A}}}_{a}\in{{\mathbb{K}}}$ by Corollary 3.6.
@@
    If $a>0$ then ${\boldsymbol{\mathit{A}}}_{a}\in{{\mathbb{K}}}_{SI}^{+}$ and thus ${\boldsymbol{\mathit{A}}}\in V_{I_{1},J_{1}}$. Consequently ${\boldsymbol{\mathit{A}}}\in{{\mathbb{K}}}$.
@@
    If $a<0$ then ${\boldsymbol{\mathit{A}}}_{a}\in{{\mathbb{K}}}_{SI}^{-}$ and thus ${\boldsymbol{\mathit{A}}}\in V^{d}_{I_{2},J_{2}}$. This implies ${\boldsymbol{\mathit{A}}}\in{{\mathbb{K}}}$.
This proves that the equations from $S_{I_{1},J_{1}}\cup S_{I_{2},J_{2}}^{d}$ axiomatize $\mathbb{K}$}}. ∎
6 Applying Mundici functor
In the previous sections, we developed a self-contained theory using the language of Abelian $\ell$-groups. In this section we will establish several connections between theory of MV-algebras and theory of pointed Abelian $\ell$-groups. We will show that the semantical description of subvarieties of $\mathrm{pAb}$$^{+}$ is mutually equivalent with the Komori classification of subvarieties of MV-algebras. We will conclude the section by deriving a complete classification of quasivarieties of $\mathrm{pAb}$} generated by totally ordered elements from similar result about MV-algebras [14].
Let us recall that MV-algebras can be understood as bounded intervals of Abelian $\ell$-groups. For us, MV-algebra will be a structure of the following signature $\langle{\oplus,\otimes,\lor,\land,\neg,0,1}\rangle$. For the basics of MV-algebra we recommend the reader to check any classical bibliography e.g. [7].
We shall note that we are adding the symbols $\lor,\land$ and $\otimes$ into our language purely for our convenience, it is well-known that $\lor,\land$ and $\otimes$ are term definable using the remaining operations.
The main tool we use in this section is Mundici functor, which is a functor from the category of strongly positively pointed Abelian $\ell$-groups into the category of MV-algebras. We will denote the Mundici functor by $\Gamma$. Here, we provide the definition of $\Gamma$ on objects.
@@ 6 Applying Mundici functor
Let ${\boldsymbol{\mathit{A}}}_{u}=\langle{A,+,-,\lor,\land,0,u}\rangle$ be a positively pointed Abelian $\ell$-group and $u\geq 0$. Define MV-algebra $\Gamma({\boldsymbol{\mathit{A}}}_{u})=\langle{[0,u],\oplus,\odot,\lor,\land,\neg,0,u}\rangle$, where the operations are defined as follows:
$a\oplus b=(a+b)\land u$
$a\otimes b=0\lor(a+b-u)$
$\neg a=u-a$
Similarly, for a homomorphism $f:{\boldsymbol{\mathit{A}}}_{u}\rightarrow{\boldsymbol{\mathit{B}}}_{v}$ we define $\Gamma(f):\Gamma({\boldsymbol{\mathit{A}}}_{u})\rightarrow\Gamma({\boldsymbol{\mathit{B}}}_{v})$ as $\Gamma(f)=f\restriction[0,u]$. It is well-known that $\Gamma$ is indeed a functor. Moreover, it is known [7] that $\Gamma$ is a categorical equivalence. We will denote its ”inverse” by $\Gamma^{-1}$.
It can be elementarily shown (as an exercise or see [28]) that $\Gamma$ and $\Gamma^{-1}$ preserve products, subalgebras and homomorphisms. To be more precise, we state this lemma.
@@
For any positively strongly pointed Abelian $\ell$-groups ${\boldsymbol{\mathit{A}}}_{u},{\boldsymbol{\mathit{B}}}_{v}$ the following hold:
@@
    $\Gamma(\prod_{i\in I}{\boldsymbol{\mathit{A}}}_{i})=\prod_{i\in I}\Gamma({\boldsymbol{\mathit{A}}}_{i})$,
@@
    ${\boldsymbol{\mathit{A}}}_{u}\in\mathbf{IS}({\boldsymbol{\mathit{B}}}_{v})$ iff $\Gamma({\boldsymbol{\mathit{A}}}_{u})\in\mathbf{IS}(\Gamma({\boldsymbol{\mathit{B}}}_{v}))$,
@@
    ${\boldsymbol{\mathit{A}}}_{u}\in\mathbf{H}({\boldsymbol{\mathit{B}}}_{v})$ iff $\Gamma({\boldsymbol{\mathit{A}}}_{u})\in\mathbf{H}(\Gamma({\boldsymbol{\mathit{B}}}_{v}))$.
@@
Every variety of MV-algebras is generated by

### v2

@@ 5 Axiomatization of subvarieties of $p\mathbb{AL}$
Now let us fix $p\in I\setminus J$. Since $p\notin J$ by the same argument as in the previous paragraph we obtain ${\boldsymbol{\mathit{Z}}}_{j}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0}\vDash\mathrm{(div_{p,n})}$ for each $j\in J$. By Lemma 5.1 we have ${\boldsymbol{\mathit{Z}}}_{i}\vDash\mathrm{(\text{s-rank}_{n})}$ for each $i\in I$ and thus by Lemma 5.5 we get ${\boldsymbol{\mathit{Z}}}_{i}\vDash\mathrm{(mix_{p,n})}$ and ${\boldsymbol{\mathit{Z}}}_{j}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0}\vDash\mathrm{(mix_{p,n})}$ for each $i\in I$ and $j\in J$. Thus $\mathcal{V}_{I,J}\vDash\mathrm{(mix_{p,n})}$ for each $p\in I\setminus J$.
@@
This concludes that $\mathcal{V}_{I,J}\vDash\mathrm{(eq)}$ for each $\mathrm{(eq)}\in S_{I,J}$. It remains to show $\mathcal{V}_{I,J}$ is uniquely determined by the equations from $S_{I,J}$.
@@
Let ${\boldsymbol{\mathit{A}}}_{a}\in p\mathbb{AL}^{+}_{\mathrm{FSI}}$ be a finitely generated satisfying all the equations from $S_{I,J}$. We will show ${\boldsymbol{\mathit{A}}}_{a}\in\mathcal{V}_{I,J}$. First, let us note that clearly $\mathbf{HSP}({\boldsymbol{\mathit{A}}}_{a})\neq\mathbf{HSP}({\boldsymbol{\mathit{R}}}_{1})$ since from Lemma 3.7 we know $\mathbf{HSP}({\boldsymbol{\mathit{R}}}_{1})=p\mathbb{AL}^{+}$ and $\boldsymbol{\mathit{A}}$} satisfies equations from $S_{I,J}$, which are not generally valid in $p\mathbb{AL}^{+}$. By Corollary 3.6 and Theorems 3.10 and 4.10 it is enough to consider the cases ${\boldsymbol{\mathit{A}}}_{a}\in\{{\boldsymbol{\mathit{Z}}}_{i},{\boldsymbol{\mathit{Z}}}_{j}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0}\mid i,j\in\mathbb{N}\}$.
@@
For ${\boldsymbol{\mathit{A}}}_{a}={\boldsymbol{\mathit{Z}}}_{m}$ we get $m\leq n$, by applying Lemma 5.4 using ${\boldsymbol{\mathit{Z}}}_{m}\vDash\mathrm{(rank_{n})}$. Since ${\boldsymbol{\mathit{Z}}}_{m}\vDash\mathrm{(div_{p,n})}$ for all $p\notin I$, we get by Lemma 5.2 that $m\in I$.
@@
For ${\boldsymbol{\mathit{A}}}_{a}={\boldsymbol{\mathit{Z}}}_{m}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0}$ we similarly get $m\leq n$ and $m\in I$. Since for every $p\in I\setminus J$ we have ${\boldsymbol{\mathit{Z}}}_{m}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0}\vDash\mathrm{(mix_{p,n})}$, by Corollary 5.6 we get $m\in J$.
Thus ${\boldsymbol{\mathit{A}}}_{a}\in\mathcal{V}_{I,J}$, which completes the proof. ∎
@@
The structural part of Theorem 5.7 was proved in [32]. This result can be obtained by applying the Mundici functor, Corollary 3.6, Theorem 3.10, and Komori classification, an approach we will discuss later in Section 6 and demonstrate in Theorems 6.3 and 6.6.
@@
Despite the fact that the statement of Theorem 5.7 was about proper varieties, we can also claim that any variety of $p\mathbb{AL}^{+}$ is equal to $\mathcal{V}_{I,J}$ for some sets $J\subseteq I$, since $p\mathbb{AL}^{+}=\mathcal{V}_{\mathbb{N},\mathbb{N}}$.
@@
Any non-trivial proper subvariety of $p\mathbb{AL}^{-}$ is of the form
$\mathcal{V}^{d}_{I,J}=\mathbf{HSP}({\boldsymbol{\mathit{Z}}}_{-i},{\boldsymbol{\mathit{Z}}}_{-j}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0}\mid i\in I,j\in J)$
for some finite sets $\{0\}\subseteq J\subseteq I\subsetneq\mathbb{N}$, where both $I\setminus\{0\}$ and $J\setminus\{0\}$ are closed under divisors.
Moreover, $\mathcal{V}^{d}_{I,J}$ is axiomatized by the following set $S_{I,J}^{d}$ of equations:
@@
Any non-trivial subvariety of $p\mathbb{AL}$ can be written as $\mathcal{V}=\mathcal{V}_{I_{1},J_{1}}\lor\mathcal{V}^{d}_{I_{2},J_{2}}$ for some (not necessarily finite) sets $I_{1},I_{2},J_{1},J_{2}$. Such a variety is axiomatized by the equations from $S_{I_{1},J_{1}}\cup S_{I_{2},J_{2}}^{d}$, where $S_{I_{1},J_{1}}$ and $S^{d}_{I_{2},J_{2}}$ are defined in Theorems 5.7 and 5.8.
[…]
Figure 1: The lattice of join irreducible subvarieties of $p\mathbb{AL}$.
@@
Let $\mathcal{V}$ be an arbitrary subvariety of $p\mathbb{AL}$. Using Theorem 2.6 we can derive that finitely subdirectly irreducible pointed Abelian $\ell$-groups are totally ordered. Let us denote $\mathcal{V}_{\mathrm{FSI}}^{+}=\mathcal{V}_{\mathrm{FSI}}\cap p\mathbb{AL}^{+}$ and $\mathcal{V}_{\mathrm{FSI}}^{-}=\mathcal{V}_{\mathrm{FSI}}\cap p\mathbb{AL}^{-}$. Thus we have
$\mathbf{HSP}(\mathcal{V}_{\mathrm{FSI}}^{+}\cup\mathcal{V}_{\mathrm{FSI}}^{-})=\mathbf{HSP}(\mathcal{V}_{\mathrm{FSI}})=\mathcal{V}.$
This shows that $\mathcal{V}$ is the join of $\mathbf{HSP}(\mathcal{V}_{\mathrm{FSI}}^{+})$ and $\mathbf{HSP}(\mathcal{V}_{\mathrm{FSI}}^{-})$. By Theorem 5.7 $\mathbf{HSP}(\mathcal{V}_{\mathrm{FSI}}^{+})=\mathcal{V}_{I_{1},J_{1}}$ for some sets $I_{1},J_{1}$ and by Theorem 5.8 $\mathbf{HSP}(\mathcal{V}_{\mathrm{FSI}}^{-})=\mathcal{V}^{d}_{I_{2},J_{2}}$ for some sets $I_{2},J_{2}$. Thus we have $\mathcal{V}=\mathcal{V}_{I_{1},J_{1}}\lor\mathcal{V}^{d}_{I_{2},J_{2}}$.
We show that the equations from $S_{I_{1},J_{1}}\cup S_{I_{2},J_{2}}^{d}$ axiomatize the variety $\mathcal{V}$. First all the equations from $S_{I_{1},J_{1}}$ are valid in all negatively pointed Abelian $\ell$-groups by Lemmas 5.2, 5.4 and 5.5. Similarly, all the equations from $S^{d}_{I_{2},J_{2}}$ are valid in all positively pointed Abelian $\ell$-groups. Since all the equations from $S_{I_{1},J_{1}}$ are valid in $\mathcal{V}_{I_{1},J_{1}}$ and all the equations from $S^{d}_{I_{2},J_{2}}$ are valid in $\mathcal{V}^{d}_{I_{2},J_{2}}$, we conclude that all equations from $S_{I_{1},J_{1}}\cup S_{I_{2},J_{2}}^{d}$ are valid in $\mathcal{V}$.
Now let us take arbitrary ${\boldsymbol{\mathit{A}}}_{a}\in p\mathbb{AL}_{\mathrm{FSI}}$ satisfying all the equations from $S_{I_{1},J_{1}}\cup S_{I_{2},J_{2}}^{d}$. By Theorem 2.6 ${\boldsymbol{\mathit{A}}}_{a}$ is totally ordered. We distinguish three cases:
@@
    If $a=0$ then ${\boldsymbol{\mathit{A}}}_{a}\in\mathcal{V}$ by Corollary 3.6.
@@
    If $a>0$ then ${\boldsymbol{\mathit{A}}}_{a}\in\mathcal{V}_{\mathrm{FSI}}^{+}$ and thus ${\boldsymbol{\mathit{A}}}_{a}\in\mathcal{V}_{I_{1},J_{1}}$. Consequently ${\boldsymbol{\mathit{A}}}_{a}\in\mathcal{V}$.
@@
    If $a<0$ then ${\boldsymbol{\mathit{A}}}_{a}\in\mathcal{V}_{\mathrm{FSI}}^{-}$ and thus ${\boldsymbol{\mathit{A}}}_{a}\in\mathcal{V}^{d}_{I_{2},J_{2}}$. This implies ${\boldsymbol{\mathit{A}}}_{a}\in\mathcal{V}$.
This proves that the equations from $S_{I_{1},J_{1}}\cup S_{I_{2},J_{2}}^{d}$ axiomatize $\mathcal{V}$. ∎
6 Applying the Mundici functor
In the previous sections, we developed a self-contained theory using the language of pointed Abelian $\ell$-groups. In this section we will establish several connections between the theory of MV-algebras and the theory of pointed Abelian $\ell$-groups. We will show that the semantical description of subvarieties of $\mathrm{pAb}$$^{+}$ is analogous to the Komori classification of subvarieties of MV-algebras. We will conclude the section by deriving a complete classification of quasivarieties of $\mathrm{pAb}$} generated by totally ordered elements from the similar result about MV-algebras [16].
Let us recall that MV-algebras can be understood as bounded intervals of Abelian $\ell$-groups. For us, an MV-algebra will be a structure of the following signature $\langle{\oplus,\otimes,\lor,\land,\neg,0,1}\rangle$. For the basics of MV-algebras we refer the reader to any classical bibliography e.g. [8].
We shall note that we are adding the symbols $\lor,\land$ and $\otimes$ into our language purely for our convenience. It is well-known that $\lor,\land$ and $\otimes$ are term definable using the remaining operations.
The fundamental tool for working with MV-algebras is the Mundici functor, which is a categorical equivalence from the category of strongly positively pointed Abelian $\ell$-groups into the category of MV-algebras. As discussed in [32], one can generalize the Mundici functor to the whole category of positively pointed Abelian $\ell$-groups. We will denote this extended functor as $\Gamma$ and define it as follows:
@@ 6 Applying the Mundici functor
Let ${\boldsymbol{\mathit{A}}}_{u}=\langle{A,+,-,\lor,\land,0,u}\rangle$ be a positively pointed Abelian $\ell$-group and $u\geq 0$. Define MV-algebra $\Gamma({\boldsymbol{\mathit{A}}}_{u})=\langle{[0,u],\oplus,\otimes,\lor,\land,\neg,0,1}\rangle$, where the operations are defined as follows:
  1. 1.
    $a\oplus b:=(a+b)\land u$,
  2. 2.
    $a\otimes b:=0\lor(a+b-u)$,
  3. 3.
    $\neg a:=u-a$,
  4. 4.
    $1:=u.$
For positively pointed Abelian $\ell$-groups ${\boldsymbol{\mathit{A}}}_{u}$, ${\boldsymbol{\mathit{B}}}_{v}$ and a homomorphism $f:{\boldsymbol{\mathit{A}}}_{u}\rightarrow{\boldsymbol{\mathit{B}}}_{v}$ we define $\Gamma(f):\Gamma({\boldsymbol{\mathit{A}}}_{u})\rightarrow\Gamma({\boldsymbol{\mathit{B}}}_{v})$ as $\Gamma(f):=f\restriction[0,u]$.
It is well-known [8] that the Mundici functor ($\Gamma$ restricted to the subcategory of strongly positively pointed Abelian $\ell$-groups) is a categorical equivalence. By $\Gamma^{-1}$ we denote the corresponding inverse functor. In particular, we have $\Gamma(\Gamma^{-1}({\boldsymbol{\mathit{B}}}))\cong{\boldsymbol{\mathit{B}}}$ for each MV-algebra $\boldsymbol{\mathit{B}}$}. It is well known [32] that for positively pointed Abelian $\ell$-group ${\boldsymbol{\mathit{A}}}_{a}$ with $a\neq 0$ we have that $\Gamma^{-1}(\Gamma({\boldsymbol{\mathit{A}}}_{a}))$ is equal to its strongly pointed convex $\ell$-subgroup. Specially, if ${\boldsymbol{\mathit{A}}}_{a}$ is itself strongly pointed, we have $\Gamma^{-1}(\Gamma({\boldsymbol{\mathit{A}}}_{a}))\cong{\boldsymbol{\mathit{A}}}_{a}$.
@@
Let ${{\mathbb{K}}}\cup\{{\boldsymbol{\mathit{A}}}_{u}\}$ be a class of positively pointed Abelian $\ell$-groups. Then it holds ${\boldsymbol{\mathit{A}}}_{u}\in\mathbf{HSP}({{\mathbb{K}}})$ implies $\Gamma({\boldsymbol{\mathit{A}}}_{u})\in\mathbf{HSP}(\Gamma[{{\mathbb{K}}}])$.
Proof.
One has to show the following:
@@
    $\Gamma(\mathbf{P}({{\mathbb{K}}}))=\mathbf{P}(\Gamma[{{\mathbb{K}}}])$ for any class ${{\mathbb{K}}}\subseteq p\mathbb{AL}^{+}$,
@@
    ${\boldsymbol{\mathit{A}}}_{u}\in\mathbf{S}({\boldsymbol{\mathit{B}}}_{v})$ implies $\Gamma({\boldsymbol{\mathit{A}}}_{u})\in\mathbf{S}(\Gamma({\boldsymbol{\mathit{B}}}_{v}))$ for any ${\boldsymbol{\mathit{A}}}_{u},{\boldsymbol{\mathit{B}}}_{v}\in p\mathbb{AL}^{+}$,
@@
    ${\boldsymbol{\mathit{A}}}_{u}\in\mathbf{H}({\boldsymbol{\mathit{B}}}_{v})$ implies $\Gamma({\boldsymbol{\mathit{A}}}_{u})\in\mathbf{H}(\Gamma({\boldsymbol{\mathit{B}}}_{v}))$ for any ${\boldsymbol{\mathit{A}}}_{u},{\boldsymbol{\mathit{B}}}_{v}\in p\mathbb{AL}^{+}$.
For detailed discussion, see [32]. ∎
@@
Every non-trivial proper variety of MV-algebras is generated by